EASY-GOING ON-SPECTROMETER OPTIMISATION OF PHASE MODULATED HOMONUCLEAR DE-COUPLING SEQUENCES IN SOLID-STATE NMR

DENNIS L. A. G. GRIMMINCK, SURESH K. VASA, \underline{W. LEO MEERTS}, AND P. M. KENT-GENS, Radboud University, Institute for Molecules and Materials, Heyendaalseweg 135, NL-6525 AJ Nijmegen, The Netherlands.

A global optimisation scheme for phase modulated proton homonuclear decoupling sequences in solid-state NMR is presented. Phase modulations, parameterised by DUMBO Fourier coefficients, were optimized using a Covariance Matrix Adaptation Evolution Strategies algorithm^{a}. Our method, denoted EASY-GOING homonuclear decoupling, starts with featureless spectra and optimises proton-proton decoupling, during either proton or carbon signal detection. On the one hand, our solutions closely resemble (e)DUMBO^{b} for moderate sample spinning frequencies and medium radio-frequency (rf) field strengths. On the other hand, the EASY-GOING approach resulted in a superior solution, achieving significantly better resolved proton spectra at very high 680 kHz rf field strength.

aN. Hansen, and A. Ostermeier. Evol. Comput. 9 (2001) 159–195

bB. Elena, G. de Paëpe, L. Emsley. Chem. Phys. Lett. 398 (2004) 532–538